OpenStudy (anonymous):

Choose the correct simplification of 3^8/3^3

OpenStudy (anonymous):

\[\frac{3^8}{3^3}=\frac{3 \times 3\times 3\times 3\times 3\times 3\times 3 \times 3}{3 \times 3 \times 3}\]Now we can cancel, is that ok?

OpenStudy (anonymous):

i got 1/15 am i right?

OpenStudy (anonymous):

\[\frac{3\times 3\times 3\times 3\times 3\times \cancel{ 3\times 3\times 3}}{\cancel{3 \times 3 \times 3}}= 3\times 3\times 3\times 3\times 3\]Does that make sense?

OpenStudy (anonymous):

so its 3^4

OpenStudy (anonymous):

@hea

OpenStudy (anonymous):

How many 3's do you count on the right hands side of my equation?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

so 3^5

OpenStudy (anonymous):

yep!